OpenStudy (anonymous):

solve the system of equations 9x-3y=36 7x-3y=45

OpenStudy (texaschic101):

multiply the 1st equation by -1 and then use the elimination method. Can you do that ?

OpenStudy (anonymous):

-9x+3y=-36 7x-3y=45 -2x=8 x=-4

OpenStudy (texaschic101):

you found the x....now sub -4 in for x in either of the equations to find y

OpenStudy (texaschic101):

wait......x is wrong...try that again

OpenStudy (anonymous):

-3y = 36 - 9x 3y = 9x - 36 y = 3x - 12 7x - 3(3x - 12) = 45 7x - 9x + 36 = 45 -2x = 9 x = -9/2 y = 3x - 12 y = 3(-9/2) - 12 y = (-27 - 24)/2 y = -51/2

OpenStudy (anonymous):

isn't the answer (9,6)

OpenStudy (texaschic101):

no.....its gonna be a fraction....huiyingtanyayap just did it for you using the substitution method

OpenStudy (texaschic101):

let me check her answers...hold on

OpenStudy (texaschic101):

9x - 3y = 36 9(-9/2) - 3(-51/2) = 36 -81/2 + 153/2 = 36 -72/2 = 36 -36 = 36... no those answers are not correct.... Let me do it....hold on

OpenStudy (texaschic101):

9x - 3y = 36 -3y = 36 - 9x (multiply by -1) 3y = 9x - 36 y = 3x - 12 7x - 3y = 45 7x - 3(3x - 12) = 45 7x - 9x + 36 = 45 -2x = 45 - 36 -2x = 9 x = -9/2 9x - 3y = 36 9(-9/2) - 3y = 36 -81/2 -3y = 36 -81/2 - 36 = 3y -81/2 - 72/2 = 3y (multiply the equation by 2) -81 - 72 = 6y -153 = 6y -153/6 = y check... 7x - 3y = 45 7(-9/2) - 3(-153/6) = 45 -63/2 + 459/6 = 45 (multiply the equation by 6) -189 + 459 = 270 270 = 270 (correct) x = -9/2 y = -153/6